Database last updated on Sun, 01 Feb 2026 17:40:48 +0000 by localhost
Instrument constant changes for the week before 2023263 (2 sec, 0 rows)
HG burned out? (AP results when HG is ON) (0 sec, 17 rows)
Measurements and tests summary.
_jday
_loc
_serial
_skc
aode
ds
dt
ed
fz
hg
rs
sl
uv
zs
263
GOOSE BAY
#018
ecsc,Next schedule will be ecsc (now ecsc ZS= 93 ZS= 93.033)
1
1
1
2
1
19
1
6
55
75
(1 sec)
Dead time (0 sec)
File name
Count
Serial
Expected (+/-5)
Measured
30-day history
B26323.018
1
#018
35
35